# Adaptive Coordinators and Prompts on Heterogeneous Graphs for Cross-Domain Recommendations

### v1

@@ 3.2.4. Multi-domain Pre-training
In order to demonstrate the wide versatility of our framework, we subsequently adopt other state-of-the-art graph pre-training methods like BGRL and COSTA, integrating them into our multi-domain pre-training framework, in Section 4.4. COSTA builds upon GRACE by incorporating the novel Covariance-Preserving Feature Augmentation technique, which effectively tackles the bias challenges present in the original augmentation strategies. BGRL introduces the concept of Bootstrapping Latent Loss, which is designed to align the node representations produced by the online and target encoders with different updating strategies.
@@ 3.3.2. Training Objective
where the second term is the $2$-regularization for potential overfitting issues. $(u,i,j)$ from batch $\mathcal{B}$ is a triple including user $u$ and item $i,j$. The interaction between user $u$ and item $i$ represents a positive pair, while the absence of interaction between user $u$ and item $j$ signifies a negative pair. $\hat{y}_{ui}=\boldsymbol{o}_{u}\cdot\boldsymbol{o}_{i}$ and $\hat{y}_{uj}=\boldsymbol{o}_{u}\cdot\boldsymbol{o}_{j}$ are the predicted value of interaction occurring between user $u$ and item $i,j$. $\Theta$ denotes the entire ensemble of trainable parameters that the model employs and $\lambda$ controls the regularization degree.
@@ 3.4. Why It Works?
Theoretical Support: Our proposed framework shares a common theoretical bedrock with graph prompts designed for knowledge transfer. The primary difference is that while graph prompts are generally targeted at downstream tasks, we utilize the term ”graph coordinators” for the pre-training phase. Studies (Wang et al., 2024) have shown that for a given graph $\mathcal{G}$, we can find some potential graph transformation $T(\cdot)$ to translate the original graph as a new one like $T(\mathcal{G})$ such that the pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ can approximate to an optimal function $C$ for the target domain task, which can be described as:
@@
It can be proven that graph prompts can be treated as a learnable method to find such data transformation (Fang et al., 2023; Sun et al., 2023a). Specifically, given a pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ with fixed parameters $\theta^{*}$ and learnable graph prompt $P(\cdot)$, it can be shown that:
@@
where $O_{fp}$ is the bounded error associated with $f_{\theta^{*}}(\cdot)$ and $P(\cdot)$. In other words, graph prompts have the capability to be equivalent to certain graph transformation operations that enable the transformed graph, once processed by the pre-trained graph encoder, to produce optimal embeddings. In this paper, our heterogeneous adaptive graph coordinators can be treated as a variant of graph prompts, which is used in the pre-training stage, and thus share the same foundation for data operations. Based on this idea, we give a data-centric solution to heal the gap among multiple source domains and the target domain such that the performance ceiling of knowledge transfer can be further improved as mentioned in our motivation (Section 2).
Compatibility Analysis: Our proposed HAGO is model-agnostic because it extends the graph structure without regarding to message passing and aggregation within the graph, making it compatible with various graph-based backbone networks. Graph Prompting can also be seamlessly integrated with various graph pre-training algorithms. In our subsequent experiments, comprehensive compatibility analysis demonstrates the broad compatibility and wide effectiveness of HAGO with different backbone networks and pre-training methods.
@@ 4.1. Experiment Settings
Single-domain Method Cross-domain Method
Data Preparation and Setup: We conduct comprehensive experiments on two popular public recommendation datasets: Amazon dataset (Ni et al., 2019) [^2] http://jmcauley.ucsd.edu/data/amazon is gathered from relevant product evaluations on Amazon.com, which are extensively utilized to assess the performance of cross-domain recommender systems. For our study, we have chosen four categories: “Books”, “Electronics”, “Movies and TV”, and “Toys and Games”; Douban dataset (Zhao et al., 2021) [^3] https://recbole.s3-accelerate.amazonaws.com/CrossDomain/Douban.zip is collected from Douban website, which is a large-scale platform for rating and reviewing movies, TV shows, books, and music. This dataset includes three domain datasets: “Douban-Movie”, “Douban-Book”, and “Douban-Music”. In our experiments, we adopt the domain with the least interactions in each dataset as the target domain, namely “Douban-Book” and “Amazon-Toy”, with the other domains serving as the source domains. More detailed statistics of the datasets are given in Table 1.
Evaluation Metrics: We assess the performance of our method by four widely used metrics in recommendations: Recall@10, HR@10 (Hit Ratio), NDCG@10 (Normalized Discounted Cumulative Gain), and MRR (Mean Reciprocal Rank) for top-10 recommendation. Kindly note that we fully rank the entire items within the datasets for calculating the above top-10 recommendation performance rather than evaluate these metrics by pre-sampling a small group of the total items. This is closer to the real industrial applications and raises more challenging tasks for the compared models.
Baselines: We compare our method with NINE state-of-the-art methods, which can be grouped into two categories: (1) Single-domain Recommendation Methods withNGCF (Wang et al., 2019), LightGCN (He et al., 2020), and XSimGCL (Yu et al., 2024). (2) Cross-domain Recommendation Methods with CMF (Singh and Gordon, 2008), EMCDR (Man et al., 2017), DeepAPF (Yan et al., 2019), SSCDR (Kang et al., 2019), BiTGCF (Liu et al., 2020), and CUT (Li et al., 2024a). The details of these methods are present in Appendix A.2.
@@ 4.2. Overall Performance
  • Our proposed HAGO consistently achieves the best performance across all four metrics on two multi-domain datasets, demonstrating its superior effectiveness. It gains 1.20% and 1.28% improvements on Recall@10, 1.44% and 1.67% on HR@10, 1.32% and 1.64% on NDCG@10, and 0.59% and 2.57% on MRR against the runner-up on Douban and Amazon datasets. HAGO owes its success to employing graph coordinators that seamlessly integrate multi-domain graphs into a cohesive framework for multi-domain collaborative pre-training. These graph coordinators adaptively connect with corresponding domain graphs, dynamically adjusting the information communication within and across domains, achieving information enhancement effectively while mitigating negative transfer. Moreover, it adopts graph prompts to efficiently transfer valuable knowledge from multiple source domains to the target domain, thereby boosting the recommendation performance.
@@ 4.4. Compatibility Analysis with Different Backbones and Pre-training Strategies
LightGCN 0.1627 0.1902 0.1141 0.1074
HAGO(LightGCN) 0.1662 0.1950 0.1177 0.1116
Our proposed HAGO is not confined to a specific model but is designed as a flexible architecture that can be effortlessly combined with a variety of backbone networks serving as graph encoders, as well as with diverse graph pre-training algorithms. To demonstrate the adaptability of HAGO, we conduct comparative experiments by integrating it with other notable graph neural networks, including GCN and GAT. Furthermore, we investigate the incorporation of other self-supervised learning strategies for pre-training the graph, such as BGRL and COSTA.
@@ 4.5. Impact Analysis of Coordinator Number
Observing the results, it is evident that the optimal performance of HAGO is reached when the number of graph coordinators is moderate. When there are too few coordinators, they are inadequate to handle the demanding task of transferring knowledge both within and across domains efficiently. As the number of coordinators increases, the marginal benefits to model performance from the additional coordinators diminish, and there may even be a decline in performance, potentially due to excessive knowledge transfer.
@@ 4.6. Visualization Analysis
To investigate whether the HAGO framework we propose can effectively align the representation spaces and facilitate cross-domain knowledge transfer among multiple domains, we conduct a visualization analysis of the pre-trained embeddings for item nodes in each domain, following the analytical methods in (Yu et al., 2022). We randomly sample 2000 items from each of the three domains in the Douban dataset (Douban-Music, Douban-Movie, and Douban-Book). Subsequently, we utilize t-SNE (Van der Maaten and Hinton, 2008) to perform dimensionality reduction to map these item embeddings to two-dimensional feature vectors in the $\mathbb{R}^{2}$ and normalize them onto the unit circle. Employing Gaussian kernel estimation (Botev et al., 2010), we plot the feature distribution and corresponding angle distribution of the item samples from each domain as they are mapped onto the unit circle.
@@
Based on this approach, we perform visualizations for the isolated pre-training without graph coordinators (w/o GO), as well as the proposed HAGO, as shown in Figure 4. From the visualization results, it is observable that the item embeddings learned by the model without GO exhibit a significant clustering effect, with the embeddings from the three domains converging in three distinct directions on the unit circle, and their angle distributions are sharper. It indicates that isolated pre-training without graph coordinators fails to capture the connections between nodes from different domains, resulting in a disjointed embedding space across domains. In contrast, for HAGO, it is evident that the feature distributions corresponding to each domain are more uniformly distributed and interweave with one another. This not only suggests that HAGO can capture higher-quality embeddings that are more evenly spread throughout the embedding space, but also that the interweaving distribution of embeddings from different domains indicates HAGO’s capability to capture cross-domain associations among items. This observation underscores the effectiveness of HAGO in enhancing the expressiveness of embeddings by promoting a more cohesive structure within multi-domain graphs, thereby facilitating knowledge transfer across domains. We give more additional case studies in Appendix A.3.
@@ 5. Conclusions
Our work delves into the multi-domain representation alignment and the challenging issue of negative transfer in cross-domain recommendation tasks. To tackle these intricate challenges, we introduce the HAGO framework, designed to integrate multi-domain graphs into a cohesive structure. Within HAGO, graph coordinators dynamically connect with corresponding graphs based on adaptive weights, thereby regulating intra-domain and inter-domain information transfer to mitigate the negative transfer problem. We conduct comprehensive experiments to demonstrate the effectiveness of HAGO and vividly discuss its role in representation alignment and the facilitation of cross-domain transfer through visual analysis and case studies.
Appendix A Appendix
The Appendix is organized as follows: We supplement a more comprehensive review of related work in Appendix A.1, then give a more detailed introduction to our compared baselines in Appendix A.2. In Appendix A.3, we extend our experimental discussion with additional case studies. We open our source code in Appendix A.4.
A.1. Related Work
A.1.1. Self-Supervised Graph Learning
Self-supervised graph learning is a machine learning algorithm that trains graph models using supervision signals generated from the graph data itself, without relying on additional labels (Sun et al., 2023b). Specifically, self-supervised graph learning algorithms can be categorized into generative, predictive, and contrastive approaches  (Liu et al., 2023; Wu et al., 2023). Generative methods typically rely on an encoder-decoder structure to capture generalized latent representations through tasks like feature or structure reconstruction, such as MGAE (Wang et al., 2017) and GALA (Park et al., 2019) for feature reconstruction, and GAE (Kipf and Welling, 2016) and GraphMAE (Hou et al., 2022, 2023) for structure reconstruction. Predictive methods use pseudo-labels generated from the data itself as supervision signals. For instance, S2GRL (Peng et al., 2020) and PairwiseDistance (Jin et al., 2020) use the prediction of the shortest path between two nodes, while M3S (Sun et al., 2020) and SHGP (Yang et al., 2022) generate pseudo-labels through clustering. Contrastive methods, which are the most widely studied, learn representations by constraining positive samples to be closer and pushing negative samples apart. These methods typically augment graph data by adding or removing edges, dropping nodes, feature transformations, or dropping features to generate a series of positive samples, then utilize loss functions such as InfoNCE  (Zhu et al., 2020; You et al., 2020; Xia et al., 2022; Zhu et al., 2022; Zhang et al., 2022), Jensen-Shannon Divergence  (Hassani and Ahmadi, 2020; Sun et al., 2021), or Bootstrapping Latent Loss  (Thakoor et al., 2021) to constrain the distance between positive and negative samples.
A.1.2. Cross-domain Recommendation
Cross-domain recommendation refers to the recommendation algorithms that enhance recommendations in one domain by transferring knowledge from other domains. These algorithms can be categorized into two types: those without overlapping users/items and those with overlapping users/ items. In non-overlapping cases, research mainly focuses on transferring or aligning semantic information across domains to facilitate knowledge transfer between the source and target domains  (Li et al., 2009; Liu et al., 2022). For overlapping cases, studies leverage shared users and/or items to transfer knowledge from the source domains to the target domain. Early attempts utilize shared representations (such as CMF (Singh and Gordon, 2008)) or concatenated representations from different domains (such as DTCDR (Zhu et al., 2019)) to make predictions, learning from both the source and target domains. Later researchers explore building cross-domain connections between neural network models to enhance knowledge transfer, such as CoNet (Hu et al., 2018) and BiTGCF (Liu et al., 2020). EMCDR (Man et al., 2017) is the first to propose a mapping function to transfer knowledge at the latent representation level, which is further expanded by models like CATN (Zhao et al., 2020) and SSCDR (Kang et al., 2019). More recent studies, such as CUT (Li et al., 2024a), have begun exploring the utilization of contrastive learning and other self-supervised techniques to transfer useful knowledge from the source to the target domain, improving recommendation accuracy.
A.2. Details of Baselines
In order to thoroughly assess the effectiveness of our proposed HAGO, we conduct a comprehensive comparison with nine state-of-the-art cross-domain baseline models:
  • Single-domain Recommendation:
    • NGCF (Wang et al., 2019) represents an early approach to using graph models to aggregate user and item node representations for collaborative filtering.
    • LightGCN (He et al., 2020) is a simplified and effective graph convolutional network model that achieves high performance.
    • XSimGCL (Yu et al., 2024) adopts contrastive learning to enhance representation learning on LightGCN.
  • Cross-domain Recommendation:
    • CMF (Singh and Gordon, 2008) is a classic cross-domain recommendation algorithm that employs a weighted loss to enable MF to jointly learn from both the source and target domains.
    • EMCDR (Man et al., 2017) introduces an embedding and mapping framework that involves training a mapping function to transfer source user embeddings into the target domain.
    • DeepAPF (Yan et al., 2019) leverages the attention network to capture both shared interests across sites and distinct interests within each site, capturing a diverse range of user preferences.
    • SSCDR (Kang et al., 2019) introduces a cross-domain mapping to transfer user-item distance information based on overlapping users across domains to address the cold-start issue for new users.
    • BiTGCF (Liu et al., 2020) improves the feature propagation to aggregate high-order connectivity in user-item graphs and transfer the knowledge by using the overlapping user as the bridge.
    • CUT (Li et al., 2024a) utilizes the target domain user similarity to filter collaborative information from the source domain during knowledge transfer across domains.
A.3. Additional Case Study
Specifically, we select the top-5 popular books that appear in the Douban-Book domain. For each selected book, we calculate its cosine similarity with all movies in the Douban-Movie domain based on the embeddings learned by HAGO, and select the corresponding top-3 relevant movies. The resulting top-5 popular books and their corresponding top-3 relevant movies for each book are presented in Table A.3.
Table 6. Case Study: Top-5 popular books in the Douban dataset, along with top-3 relevant movies for each book thought by HAGO.
Douban-Book Douban-Movie
Item ID Item Name Item Tag Item ID Item Name Item Tag
Japanese Literature, Romance, Novel, Japan, Youth, Foreign Literature, Literature, Coming of Age. 1292220 Love Letter Unrequited Love, Aesthetic, Drama, First Love, Classic, Touching, Youth, Warmth, Romance, Literary, Japan.
1046265 Norwegian Wood 1293318 Grave of the Fireflies Family, Japanese Animation, Anime, Drama, Heartwarming, Japan, War, Animation.
1291560 My Neighbor Totoro Family, Childhood, Classic, Heartwarming, Growth, Healing, Warmth, Adventure, Fantasy, Fairy Tale, Japan, Animation.
Fairy Tales, Classic, Coming of Age, Philosophy, Foreign Literature, France, Literature, Novel. 1293839 Roman Holiday Drama, American Film, Romance, Love, America, Artistic, Film, Comedy.
1084336 The Little Prince /Le Petit Prince 1291546 Farewell My Concubine Cultural Revolution, Drama, Classic, Life, History, Romance, Humanity, Homosexuality, Literary, Artistic.
1292722 Titanic Touching, Drama, Romantic, Classic, Heartwarming, Powerful, Blockbuster, American, Disaster, Love.
Chinese Literature, Classic, Novel, Marriage, Modern Literature, Literature, Life, Romance. 1293964 Waterloo Bridge Drama, Classic, American Movies, War, American, Black and White, Romance, Movies.
1008145 Fortress Besieged /Wéichéng 1301753 The Lion King Inspirational, Family, Childhood, Classic, Coming of Age, American, Warmth, Adventure, Song and Dance, Animation.
1292271 The Graduate Drama, Classic, Life, Coming of Age, American, Love, Humanity, Youth, Comedy, Ethics.
The Unbearable Lightness of Being /Nesnesitelná lehkost bytí Foreign Literature, Novel, Philosophy, Czech, Literature, Life. 1293071 Wild Strawberries /Smultronstället European Movie, Drama, Family, Classic, Black and White, Romance, Europe, Sweden.
1017143 1439205 Panda Kopanda Japanese Animation, Ghibli, Japan, Animation.
1299080 Rear Window Crime, Thriller, Classic, American, Suspense, American Movies.
Mystery, Detective, Religious, Novel, Foreign Literature, American. 1916936 Death Note Japan, animation, anime, deduction, suspense, Japanese animation, Japanese anime, Cartoon.
1040771 The Da Vinci Code 1292262 Once Upon a Time in America Crime, Drama, Gangster, American Film, Classic, Life, Coming of Age, American.
1293544 The Silence of the Lambs Crime, Thriller, Drama, American Film, Classic, Human Nature, Psychology, American, Mystery, Horror.
For a fair comparison with other state-of-the-art baseline methods, we employ the most typical ID-based recommendation paradigm (i.e., only use user and item IDs), without using other features such as item names and tags beyond our research focus. However, as observed from Table A.3, HAGO captures many intriguing associations without the use of these semantic features: For example, the movies most closely associated with the novel “Norwegian Wood” are all Japanese films with emotional themes; those related to “The Little Prince” are some of the most classic Chinese and international films talking about love; and those linked to the novel “The Da Vinci Code” are all crime or mystery movies.
A.4. Source Code
This paper is currently under review in a conference, we have submitted the source code to the venue. The source link contains the venue information so we will release it after acceptance.

### v2

@@ 3.2.4. Multi-domain Pre-training
In order to demonstrate the wide versatility of our framework, we subsequently adopt other state-of-the-art graph pre-training methods like BGRL (Thakoor et al., 2021) and COSTA (Zhang et al., 2022), integrating them into our multi-domain pre-training framework, in Section 4.4. COSTA builds upon GRACE by incorporating the novel Covariance-Preserving Feature Augmentation technique, which effectively tackles the bias challenges present in the original augmentation strategies. BGRL introduces the concept of Bootstrapping Latent Loss, which is designed to align the node representations produced by the online and target encoders with different updating strategies.
@@ 3.3.2. Training Objective
where the second term is the $L2$-regularization to prevent overfitting. $(u,i,j)$ from batch $\mathcal{B}$ is a triple including user $u$ and item $i,j$. The interaction between user $u$ and item $i$ represents a positive pair, while the absence of interaction between user $u$ and item $j$ signifies a negative pair. $\hat{y}_{ui}=\boldsymbol{o}_{u}\cdot\boldsymbol{o}_{i}$ and $\hat{y}_{uj}=\boldsymbol{o}_{u}\cdot\boldsymbol{o}_{j}$ are the predicted value of interaction occurring between user $u$ and item $i,j$. $\Theta$ denotes the entire ensemble of trainable parameters that the model employs and $\lambda$ controls the regularization degree.
@@ 3.4. Why It Works?
Theoretical Support: Our proposed framework shares a common theoretical bedrock with graph prompts designed for knowledge transfer. We draw inspiration from the “Prompt as Graph” paradigm (Zi et al., 2024; Sun et al., 2023b), which introduces multiple additional prompt nodes with the inner structure, which correspond to learnable embeddings, into the original graph and connects them through specific structures (Sun et al., 2023a; Huang et al., 2023). The primary difference is that while graph prompts are generally targeted at downstream tasks, we utilize the term “graph coordinators” for the pre-training phase. Previous studies (Fang et al., 2023; Sun et al., 2023a) have demonstrated that graph prompts can emulate any graph operations, such as node/edge addition and deletion. Recent studies (Wang et al., 2024) have shown that for a given graph $\mathcal{G}$, we can find some potential graph transformation $T(\cdot)$ to translate the original graph as a new one like $T(\mathcal{G})$ such that the pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ can approximate to an optimal function $C$ for the target domain task, which can be described as:
@@
In other words, graph prompts can be treated as a learnable method to find such data transformation (Fang et al., 2023; Sun et al., 2023a) to approximate optimal in the target task (Wang et al., 2024). Specifically, given a pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ with fixed parameters $\theta^{*}$ and learnable graph prompt $P(\cdot)$, it can be shown that:
@@
where $O_{fp}$ is the bounded error associated with $f_{\theta^{*}}(\cdot)$ and $P(\cdot)$. This implies that graph prompts have the capability to be equivalent to certain graph transformation operations that enable the transformed graph, once processed by the pre-trained graph encoder, to produce optimal embeddings. In this paper, our heterogeneous adaptive graph coordinators can be treated as a variant of graph prompts, which are used in the pre-training stage, and thus share the same foundation for data operations. Based on this idea, we give a data-centric solution to heal the gap among multiple source domains and the target domain such that the performance ceiling of knowledge transfer can be further improved as mentioned in our motivation (Section 2).
Compatibility Analysis: Our proposed HAGO is model-agnostic because it extends the graph structure without regarding to message passing and aggregation within the graph, making it compatible with various graph-based backbone networks. It can also be seamlessly integrated with various graph pre-training algorithms. In our subsequent experiments, comprehensive compatibility analysis demonstrates the broad compatibility and wide effectiveness of HAGO with different backbones and pre-training methods.
@@
Single-domain Method Cross-domain Method
@@ 4.1. Experiment Settings
Data Preparation and Setup: We conduct comprehensive experiments on two popular public datasets: Amazon dataset (Ni et al., 2019) [^1] http://jmcauley.ucsd.edu/data/amazon is gathered from relevant product evaluations on Amazon.com. For our study, we have chosen four categories: “Books”, “Electronics”, “Movies and TV”, and “Toys and Games”; Douban dataset (Zhao et al., 2021) [^2] https://recbole.s3-accelerate.amazonaws.com/CrossDomain/Douban.zip is collected from Douban website, which is a large-scale platform for rating and reviewing movies, TV shows, books, and music. This dataset includes three domain datasets: “Douban-Movie”, “Douban-Book”, and “Douban-Music”. In our experiments, we adopt the domain with the least interactions in each dataset as the target domain, namely “Douban-Book” and “Amazon-Toy”, with the other domains serving as the source domains. More detailed statistics of the datasets are given in Table 2.
Evaluation Metrics: We assess the performance of our method by four widely used metrics in recommendations: Recall@10, HR@10 (Hit Ratio), NDCG@10 (Normalized Discounted Cumulative Gain), and MRR (Mean Reciprocal Rank) for top-10 recommendation. Kindly note that we fully rank the entire items within the datasets for calculating the above top-10 recommendation performance rather than evaluating these metrics by pre-sampling a small group of the total items. This is closer to the real industrial applications and raises more challenging tasks for the compared models.
Baselines: We compare our method with NINE state-of-the-art (SOTA) methods, which can be grouped into two categories: (1) Single-domain Recommendation Methods with NGCF (Wang et al., 2019), LightGCN (He et al., 2020), and XSimGCL (Yu et al., 2024). (2) Cross-domain Recommendation Methods with CMF (Singh and Gordon, 2008), EMCDR (Man et al., 2017), DeepAPF (Yan et al., 2019), SSCDR (Kang et al., 2019), BiTGCF (Liu et al., 2020), and CUT (Li et al., 2024a).
@@ 4.2. Overall Performance
  • Our proposed HAGO consistently achieves the best performance across all four metrics on two multi-domain datasets, demonstrating its superior effectiveness. HAGO owes its success to employing graph coordinators that seamlessly integrate multi-domain graphs into a cohesive framework for multi-domain collaborative pre-training. These graph coordinators adaptively connect with corresponding domain graphs, dynamically adjusting the information communication within and across domains, achieving information enhancement effectively while mitigating negative transfer. Moreover, it adopts graph prompts to efficiently transfer valuable knowledge from multiple source domains to the target domain, thereby boosting the recommendation performance.
@@ 4.4. Compatibility Analysis with Different Backbones and Pre-training Strategies
Our proposed HAGO is not limited to a specific model but is designed as a flexible architecture that can be effortlessly combined with a variety of backbone networks serving as graph encoders, as well as with diverse graph pre-training algorithms. To demonstrate the adaptability of HAGO, we conduct comparative experiments by integrating it with other notable graph neural networks, including GCN (Kipf and Welling, 2017) and GAT (Velickovic et al., 2018). Furthermore, we investigate the incorporation of other self-supervised learning strategies for pre-training the graph, such as BGRL (Thakoor et al., 2021) and COSTA (Zhang et al., 2022).
@@ 4.5. Impact Analysis of Coordinator Number
Observing the results, it is evident that the optimal performance of HAGO is reached when the number of coordinators is moderate. When there are too few coordinators, they are inadequate to handle the demanding task of transferring knowledge both within and across domains efficiently. As the number of coordinators increases, the marginal benefits to model performance from the additional coordinators diminish, and there may even be a decline in performance, potentially due to excessive knowledge transfer.
@@ 4.6. Visualization Analysis
To investigate whether the proposed HAGO framework can effectively align the representation spaces and facilitate cross-domain knowledge transfer among multiple domains, we conduct a visualization analysis of the pre-trained embeddings for item nodes in each domain, following the analytical methods in (Yu et al., 2022). We randomly sample 2000 items from each of the three domains in the Douban dataset (Douban-Music, Douban-Movie, and Douban-Book). Subsequently, we utilize t-SNE (Van der Maaten and Hinton, 2008) to perform dimensionality reduction to map these item embeddings to two-dimensional feature vectors in the $\mathbb{R}^{2}$ and normalize them onto the unit circle. Employing Gaussian kernel estimation (Botev et al., 2010), we plot the feature distribution and corresponding angle distribution of the item samples from each domain as they are mapped onto the unit circle.
@@
Based on this approach, we perform visualizations for the isolated pre-training without graph coordinators (w/o GO), as well as the proposed HAGO, as shown in Figure 4. From the visualization results, it is observable that the item embeddings learned by the model without GO exhibit a significant clustering effect, with the embeddings from the three domains converging in three distinct directions on the unit circle, and their angle distributions are sharper. It indicates that isolated pre-training without graph coordinators fails to capture the connections between nodes from different domains, resulting in a disjointed embedding space across domains. In contrast, for HAGO, it is evident that the feature distributions corresponding to each domain are more uniformly distributed and interweave with one another. This not only suggests that HAGO can capture higher-quality embeddings that are more evenly spread throughout the embedding space, but also that the interweaving distribution of embeddings from different domains indicates HAGO’s capability to capture cross-domain associations among items. This observation underscores the effectiveness of HAGO in enhancing the expressiveness of embeddings by promoting a more cohesive structure within multi-domain graphs, thereby facilitating knowledge transfer across domains.
4.7. Case Study
Specifically, we select the most popular books under four specific themes (including romance, thriller, historical, and fantasy) that appear in the Douban-Book domain. For each selected book, we calculate its cosine similarity with all movies in the Douban-Movie domain based on the item embeddings learned by HAGO, and select the corresponding top-3 relevant movies. The resulting most popular books and their corresponding top-3 relevant movies for each book are presented in Figure 5.
Figure 5. Case Study: The most popular books under four specific themes in the Douban dataset, along with top-3 relevant movies for each book thought by HAGO.
For a fair comparison with other SOTA baseline methods, we employ the most typical ID-based recommendation paradigm (i.e., only use user and item IDs), without using other features such as item names and tags beyond our research focus. However, as observed from Figure 5, HAGO captures many intriguing associations without the use of these semantic features: For instance, the movies most closely associated with the romance novel “Norwegian Wood” are all Japanese films with emotional themes (like love, family, and friendship). In contrast, those tied to the thriller novel “The Da Vinci Code” are primarily crime or mystery genres. Moreover, users who enjoy the historical novel “Stories about Ming Dynasty” may also like some movies talking about World War II or Chinese historical stories, while movies associated with “Harry Potter and the Philosopher’s Stone” are all children’s favorites such as animations, Chinese Xianxia, and Disney adventures.
@@ 5. Conclusions
Our work delves into the multi-domain representation alignment and the challenging issue of negative transfer in cross-domain recommendation tasks. To tackle these intricate challenges, we introduce the HAGO framework, designed to integrate multi-domain graphs into a cohesive structure. We conduct comprehensive experiments to demonstrate the effectiveness of HAGO and vividly discuss its role in representation alignment and the facilitation of cross-domain transfer through visual analysis and case studies.
Acknowledgements.
This work is sponsored by Tencent WeChat Rhino-Bird Focused Research Program. Xiangguo Sun and Hong Cheng are supported by project #MMT-p2-23 of the Shun Hing Institute of Advanced Engineering, The Chinese University of Hong Kong and by grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14217622). Chengzhi Piao is supported by grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (No. 12202324).